# Floquet-engineered moiré quasicrystal patterns of ultracold Bose gases in twisted bilayer optical lattices

### v4

@@ Abstract
We investigate the formation of moiré quasicrystal patterns in Bose gases confined in twisted bilayer optical lattices via Floquet-engineered intralayer atomic interactions. Dynamical evolutions of the inverse participation ratio and the density wave amplitude reveal the stage for the emergence of moiré quasicrystal patterns, whose symmetry can be easily manipulated by the modulation frequencies and amplitudes. Reducing the frequencies and increasing the amplitudes can both facilitate lattice symmetry breaking and the subsequent emergence of rotational symmetry. Notably, a twelve-fold quasicrystal pattern emerges under specific parameters, closely resembling the moiré quasicrystal in twisted bilayer graphene. The momentum-space distributions also exhibit high rotational symmetry, which is consistent with the real-space patterns at specific evolution times. Our findings establish a new quantum platform for exploring quasicrystals and their symmetry properties in ultracold bosonic systems.
@@ II Theoretical Model
To construct spin-dependent twisted bilayer hexagonal lattices, we choose two spin states of the 87Rb atom $|1\rangle\equiv|F=1,m_{F}=0\rangle$ and $|2\rangle\equiv|F=2,m_{F}=0\rangle$ [51]. Here, $F$ and $m_{F}$ are the angular quantum number and the magnetic quantum number of the 87Rb ground state. The atoms in these two spin states are modulated by lasers with wavelengths of $\lambda_{1}=790$ nm and $\lambda_{2}=788$ nm, respectively. Three laser beams of these two wavelengths intersect at $120^{\circ}$ in the x-y plane to produce two hexagonal lattices $V_{1}$ and $V_{2}$, where each beam is linearly polarized in-plane. A hexagonal lattice potential is given by $V(\bm{r})=-V_{0}\left|\sum_{j=1}^{3}\epsilon_{j}\exp[i\bm{k}_{j}\cdot(\bm{r}-\bm{r_{0}})]\right|^{2}$, with the lattice depth $V_{0}$, the laser wave vector $k_{j}$ and the polarization $\epsilon_{j}$. These two spin-dependent lattices are twisted by $\theta/2$ each other, where the $\theta$ satisfies $\cos\theta=\frac{n^{2}+m^{2}+4mn}{2(n^{2}+m^{2}+mn)}$ with two integers $(m,n)$ [52]. The system is loaded into a harmonic trap and z-direction is tightly confined by a spin-independent potential, reducing the dynamics to quasi-2D. Meanwhile, the interlayer coupling $\Omega$ of two spin states is controlled by microwaves.
@@
where $V_{trap}=\frac{1}{2}m(\omega_{x}^{2}x^{2}+\omega_{y}^{2}y^{2}+\omega_{z}z^{2})$ is a harmonic trap and $\psi_{i}(i=1,2)$ are the wave fuctions normalized as $\int\int\left(|\psi_{i}|^{2}\right)dxdy=N$, with $N$ the total atom number. $g_{ii},g_{ij}$ characterize the intralayer and interlayer atomic interactions strength. Focus on the SU(2) symmetric interaction, we set $g_{11}=g_{22}=g_{12}=g_{0}$. In our numerics, the trapping frequency is $\omega_{x,y,z}=2\pi\times\{20,20,1000\}$ Hz, hence the system is seen as quasi-2D because of the strong potential in z-direction. The quasi-2D interaction satisfies $g_{2D}=g/{\sqrt{2\pi a_{z}}}$, where $g={4\pi\hbar^{2}a}/m$ and $a_{z}=\sqrt{{\hbar}/m\omega_{z}}$ represents the characteristic length along the z axis. Here the scattering length is $a=100a_{0}$ with $a_{0}$ the Bohr radius.
We solve the above GP equations numerically via the imaginary time evolution method to obtain the ground state. The scattering length oscillates as $a(t)=a_{0}+a_{m}\sin{\omega t}$ by modulating the magnetic field near a Feshbach resonance [53]. This modulation only changes the intralayer atomic interactions $g_{11}(t)=g_{22}(t)$, while the interlayer-atomic interactions keeps $g_{12}=g_{0}$. We use the dynamical GP equations to investigate the evolution of driven DW patterns in both real and momentum spaces.
Figure 2: IPR of the system. (a) IPR of ground DW with different lattice depth $V_{0}$ and interlayer coupling $\Omega$, where the red star denotes the parameters employed in our following Floquet engineering. (b) IPR of dynamical DW with $\omega=200$ Hz and $a_{m}=50a_{0}$.
III Inverse participation ratio
@@
$\text{IPR}=\frac{\int|\psi|^{4}d\bm{r}}{(\int|\psi|^{2}d\bm{r})^{2}},$ (2)
where we set $\int|\psi|^{2}d\bm{r}=1$. Figure 2(a) illustrates the dependence of IPR on the lattice depth $V_{0}$ and interlayer coupling $\Omega$. The IPR increases with $V_{0}$, indicating a more localized DW distribution. In contrast, it exhibits a non-monotonic dependence on $\Omega$, first decreasing and then increasing. The minimum of the IPR-$\Omega$ curve shifts toward $\Omega=1E_{r}$ as $V_{0}$ increases, where the recoil energy $E_{r}=\hbar^{2}k^{2}/2m$ serves as the energy unit. Based on this dependence, we select $\Omega=1E_{r}$ and shallow lattice depth $V_{0}=4E_{r}$ as basic parameters for our numerics (red star in Fig. 2(a)).
Additionally, we plot the IPR of the dynamical DW (Fig. 2(b)). Pronounced plateaus, increases, and subsequent decreases in the IPR are observed during specific time intervals, indicating distinct evolution stages. These stages will be quantitatively analyzed in the following section. The apparent IPR variations may signal a novel quantum dynamical phase transition.
@@ IV Moiré quasicrystal patterns
We use dynamical DW $n(\bm{r},t)=|\psi(\bm{r},t)|^{2}$ and $n(\bm{k},t)=|\psi(\bm{k},t)|^{2}$ to investigate the evolution of patterns. To quantitatively characterize the stages of DW evolution, we extract the DW amplitude [56], $A_{m}=\int\tilde{n}(\mathbf{k})d\mathbf{k}/n_{0}$ with the the Fourier transformation $\tilde{n}(\mathbf{k})=\int e^{-i\mathbf{k}\cdot\mathbf{r}}{n}(\mathbf{r})d\mathbf{r}/2\pi$. Here, $n_{0}$ is the average density before modulation. The other parameters used in our numerics are $V_{0}=4E_{r}$, $\Omega=1E_{r}$, $N=2\times 10^{4}$ and $\theta=9.43^{\circ}$ (m=3, n=4).
Figure 3: The DW amplitude and patterns with $\omega=200$ Hz and $a_{m}=50a_{0}$. (a) Four evolution stages divided by DW amplitude (blue curve) and energy (brown curve). (b) The real-space patterns at $t=10,50,64,100$ ms (top) and the corresponding momentum-space patterns amplified by logarithm (bottom). The red triangles in (a) mark the four selected temporal nodes displayed in (b).
IV.1 The real-space Moiré quasicrystal patterns
@@
Chin et al. experimentally reported the hexagonal lattice pattern formation in driven BECs, characterized by six distinct modes in momentum space [47]. In this letter, we investigate the system using the coupled GP equations, where the DW distribution in momentum space condenses to a point. Therefore, we amplify the momentum-space DW via a logarithmic function $\ln{n(\bm{k},t)}$. This amplified DW distribution similarly exhibits six distinct modes under lattice symmetry, with the modes equally spaced by $\pi/3$ in directions. In particular, each mode comprises two sub-modes, which is a direct consequence of the bilayer lattices. The angle between these sub-modes precisely matches the twist angle of the bilayer lattices (Fig. 3(b) bottom left).
Figure 4: The dynamical evolution of real-space patterns with different modulation frequencies. (a) The patterns with $a_{m}=50a_{0}$ and $\omega=200$ Hz (top), $\omega=252$ Hz (middle) and $\omega=300$ Hz (bottom). (b) The corresponding DW amplitudes of (a).
@@
The rotational symmetry of the patterns significantly depends on the modulation amplitude $a_{m}$ and frequency $\omega$. We first fix the modulation amplitude and vary the frequency. Increasing the modulation frequency breaks the rotational symmetry of the system. For example, at $a_{m}=50a_{0}$ and low $\omega$, the pattern exhibits $D_{6}$ symmetries and other distinct $D_{6}$ patterns subsequently emerge during the third stage. Each of these non-subharmonic patterns represent a new symmetrical structure. Upon increasing $\omega$ to approximately 252 Hz, the DW amplitude decreases significantly, indicating only a minor DW alteration and patterns with lattice characteristics (Fig. 4(b) blue curve). The blue curve in Fig. 4(b) is notably smoother than the other two. This occurs because this frequency corresponds to the system’s resonance frequency, leading to direct excitation without a preparation stage. For frequencies above the resonance frequency, the formation of moiré quasicrystal patterns is suppressed, while the duration of the lattice patterns is extended.
@@
IV.2 The momentum-space Moiré quasicrystal patterns
We further analyze the symmetry of the DW distribution in momentum space based on Fourier analysis $\psi(\bm{k})=\int e^{-i\mathbf{k}\cdot\mathbf{r}}{\psi}(\mathbf{r})d\mathbf{r}/2\pi$. The Fourier DW $n(\bm{k})$ also exhibits high rotational symmetry, typically forming $D_{6}$ patterns. These momentum-space patterns are consistent with real-space patterns at specific times. At $a_{m}=100a_{0}$ and $\omega=200$ Hz, the momentum-space pattern forms a hexagram nearly identical to the real-space pattern at 36.2 ms, but rotated clockwise by $60^{\circ}$. Similarly, at $a_{m}=75a_{0}$ and $\omega=200$ Hz, the patterns in both spaces match well at 42, 44, and 64.5 ms (Fig. 6(b)). At 64.5 ms in particular, both patterns exhibit a symmetrical butterfly structure. The momentum-space patterns also exhibit no subharmonicity. However, unlike the real-space patterns, some momentum-space patterns exhibit only $D_{4}$ symmetries at the center.

### v5

@@ Abstract
We investigate the formation of moiré quasicrystal patterns in Bose gasses confined in twisted bilayer optical lattices via Floquet-engineered intralayer atomic interactions. Dynamical evolutions of the total density wave amplitude exhibit the stage for the emergence of moiré quasicrystal patterns, where the pattern formation is closely associated with the momenta of collective modes excited by the weak periodic drive. Through analyzing the radial and angular density wave amplitude, we find that these new collective modes are only coupled radially and cannot be decoupled eventually. The symmetry of quasicrystal patterns can be easily manipulated by the modulation frequencies and amplitudes. Reducing the frequencies and increasing the amplitudes can both facilitate lattice symmetry breaking and the subsequent emergence of rotational symmetry. Notably, a twelve-fold quasicrystal pattern emerges under specific parameters, closely resembling the moiré quasicrystal in twisted bilayer graphene. The momentum-space distributions also exhibit high rotational symmetry, which is consistent with the real-space patterns at specific evolution times. Our findings establish a new quantum platform for exploring quasicrystals and their symmetry properties in ultracold bosonic systems.
@@ II Theoretical Model
The twisted hexagonal bilayer optical lattice of ultracold bosonic system can be realized using synthetic dimension techniques in current experiment [31, 41, 43]. Three laser beams of these two wavelengths intersect at $120^{\circ}$ in the x-y plane to produce two hexagonal lattices $V_{1}$ and $V_{2}$, where each beam is linearly polarized in-plane. A hexagonal lattice potential is given by $V(\boldsymbol{r})=-V_{0}\left|\sum_{j=1}^{3}\epsilon_{j}\exp[i\boldsymbol{k}_{j}\cdot(\boldsymbol{r}-\boldsymbol{r_{0}})]\right|^{2}$, with the lattice depth $V_{0}$, the laser wave vector $k_{j}$ and the polarization $\epsilon_{j}$. These two spin-dependent lattices are twisted by $\theta/2$ each other, where the $\theta$ satisfies $\cos\theta=\frac{n^{2}+m^{2}+4mn}{2(n^{2}+m^{2}+mn)}$ with two integers $(m,n)$ [46]. The system is loaded into a harmonic trap and z-direction is tightly confined by a spin-independent potential, reducing the dynamics to quasi-2D. Meanwhile, the interlayer coupling $\Omega$ of two spin states is controlled by microwaves.
@@
where $V_{trap}=\frac{1}{2}m(\omega_{x}^{2}x^{2}+\omega_{y}^{2}y^{2}+\omega_{z}z^{2})$ is a harmonic trap and $\psi_{i}(i=1,2)$ are the wave fuctions normalized as $\int\int\left(|\psi_{i}|^{2}\right)dxdy=N$, with $N$ the total atom number. $g_{ii},g_{ij}$ characterize the intralayer and interlayer atomic interactions strength. Focus on the SU(2) symmetric interaction in calculations, we set $g_{11}=g_{22}=g_{12}=g_{0}$ because of the similarity in scattering lengths $a_{11}$, $a_{22}$ and $a_{12}$ for the 87Rb atoms. In our numerics, the trapping frequency is $\omega_{x,y,z}=2\pi\times\{20,20,1000\}$ Hz, hence the system is seen as quasi-2D because of the strong potential in z-direction. The quasi-2D interaction satisfies $g_{2D}=g_{0}/{\sqrt{2\pi}a_{z}}$, where $g_{0}=4\pi\hbar^{2}a/m$ and $a_{z}=\sqrt{{\hbar}/m\omega_{z}}$ represents the characteristic length along the z axis. Here the scattering length for the 87Rb atoms is $a=100a_{0}$ with $a_{0}$ the Bohr radius.
The GP equations is numerically solved via the imaginary time evolution method to obtain the ground state. We make the scattering length oscillate as $a(t)=100a_{0}+a_{m}\sin{\omega t}$ by modulating the magnetic field near a Feshbach resonance [9]. This paper mainly investigates the oscillating intralayer atomic interactions $g_{11}(t)=g_{22}(t)=4\pi\hbar^{2}a(t)/(\sqrt{2\pi}a_{z}m)$, while keeping the interlayer-atomic interactions $g_{12}=g_{0}$. Then we use the dynamical GP equations to investigate the evolution of driven DW patterns in both real and momentum spaces.
Figure 2: IPR of the system. (a) IPR of ground DW with different lattice depth $V_{0}$ and interlayer coupling $\Omega$, where the red star denotes the parameters employed in our following Floquet engineering. (b) IPR and MSD
of dynamical DW with $\omega=200$ Hz and $a_{m}=50a_{0}$.
III Dynamical behavior
@@
$\text{IPR}=\frac{\int|\psi|^{4}d\boldsymbol{r}}{(\int|\psi|^{2}d\boldsymbol{r})^{2}},$ (2)
where we set $\int|\psi|^{2}d\boldsymbol{r}=1$. Figure 2(a) illustrates the dependence of IPR on the lattice depth $V_{0}$ and interlayer coupling $\Omega$. The IPR increases with $V_{0}$, indicating a more localized DW distribution. In contrast, it exhibits a non-monotonic dependence on $\Omega$, first decreasing and then increasing. The minimum of the IPR-$\Omega$ curve is around $\Omega=E_{r}$ with energy unit $E_{r}=\hbar^{2}k^{2}/2m$, indicating the best superfluidity. In addition, the GP equations is used to describe the superfluid BECs in shallow optical lattices, but $V_{0}$ must remain sufficiently large to ensure a well-defined lattice distribution. Based on this condition, we select $\Omega=E_{r}$ and $V_{0}=4E_{r}$ as basic parameters for our numerics (red star in Fig. 2(a)). The other parameters used in our numerical computations are $N=2\times 10^{4}$ and $\theta=9.43^{\circ}$ (m=3, n=4).
However, the IPR is always at the order $N^{-1}$ in the whole dynamical evolution, whose variation is too small to reflect the localization of the system. Therefore, we introduce the mean-square displacement (MSD) $\sigma(t)=[\frac{\int r^{2}|\psi|^{2}d\boldsymbol{r}}{\int|\psi|^{2}d\boldsymbol{r}}]^{\frac{1}{2}}$ [45, 1, 63] to quantitatively distinguish the different dynamical behaviors of the system at a long evolution time. $\sigma(t)$ keep a stable value at early evolution time, indicating the system remains unexcited and exhibits localized behavior. Then $\sigma(t)$ tends to linearly increase in a large evolution time, characterizing its delocalized behavior (Fig. 2(b) red curve). Notably, this linear increase comprises different segments with distinct slopes, implying four evolution stages.
@@ IV Moiré quasicrystal patterns
IV.1 Density wave amplitude
[…]
@@ IV.1 Density wave amplitude
Chin et al. experimentally reported the hexagonal lattice pattern formation in driven BECs, characterized by six distinct modes in momentum space [61]. In this letter, we investigate the system using the coupled GP equations, where the DW distribution in momentum space condenses to a point. Therefore, we amplify the momentum-space DW via a logarithmic function $\ln{n(\boldsymbol{k},t)}$. This amplified DW distribution similarly exhibits six distinct modes under lattice symmetry, with the modes equally spaced by $\pi/3$ in directions. In particular, each mode comprises two sub-modes, which is a direct consequence of the bilayer lattices. The angle between these sub-modes precisely matches the twist angle of the bilayer lattices (Fig. 3(b) bottom left).
Figure 3: The total DW amplitude and patterns with $\omega=200$ Hz and $a_{m}=50a_{0}$. (a) Four evolution stages divided by $A_{total}$
(blue curve) and energy (brown curve). (b) The real-space patterns at $t=10,50,64,100$ ms (top) and the corresponding momentum-space patterns amplified by logarithm (bottom). The red triangles in (a) mark the four selected temporal nodes displayed in (b).
@@
Figure 4: The radial and angular DW amplitudes with $\omega=200$ Hz and $a_{m}=50a_{0}$. (a) The dependence of $A_{\mathbf{k}}$ on $\mathbf{k}$ at $t=10,50,64,100$ ms, where $\mathbf{k}_{f}=\sqrt{m\omega/\hbar}$ is the the parametric resonance condition which only considers weak atomic interaction term. (b) The dynamical $A_{\mathbf{k}}$ with different momenta $\mathbf{k}$. (c) The dependence of $A_{\phi}$ on $\phi$ when $\mathbf{k}\leq\mathbf{k}_{3}$ (left) and $\mathbf{k}>\mathbf{k}_{3}$ (right) at $t=10,50,64,100$ ms.
In our calculations, the strength of the drive for one layer is $E_{g}=\left\langle|\psi_{1}|^{2}\right|g_{11}(t=0)\left||\psi_{1}|^{2}\right\rangle\approx 0.02$, which is weak compared with other energy scales ($E_{V_{1}}=\left\langle|\psi_{1}|^{2}\right|V_{1}\left||\psi_{1}|^{2}\right\rangle\approx 2.36$, $E_{\Omega}=\left\langle|\psi_{1}|^{2}\right|\Omega\left||\psi_{1}|^{2}\right\rangle\approx 0.50$). Obviously, the energies keep $E_{V_{1}}>E_{\Omega}\gg E_{g}$ in the whole dynamical evolution. Therefore, we think the formation of Moiré quasicrystal patterns is associated with momenta of collective modes that are resonantly excited by the periodic drive. To quantitatively analyze these momenta, we extract the DW amplitude, $A_{\mathbf{k}}=n_{0}^{-1}\int_{\left|\mathbf{k}\right|}\tilde{n}(\mathbf{k})d\mathbf{k}$ [14]. However, different from the Ref. [14], our system is more complex and not easy to get an analytical derivation of the parametric resonance condition from the Mathieu equation. Therefore, we numerically calculate the dependence of $A_{\mathbf{k}}$ on $\mathbf{k}$ during the dynamical evolution.
For ground state, the DW is controlled by six distinct momenta (Fig. 4a red curve), where the two main momenta ($\mathbf{k}_{1}=0.03k_{r}$, $\mathbf{k}_{2}=2k_{r}$) correspond precisely to the central mode and its surrounding six modes in the momentum space (Fig. 3(b) bottom line). A multitude of additional collective modes emerge during the pattern-formation stage ((Fig. 4a blue curve). We track the evolution of these modes throughout the dynamics and uncover several interesting regularities.
In the region of $0<\mathbf{k}<\mathbf{k}_{1}$, $A_{\mathbf{k}}$ keeps decreasing after excitation, indicating these momentum-associated collective modes play a waning influence on pattern formation. In the region of $\mathbf{k}_{1}<\mathbf{k}<\mathbf{k}_{2}$, $A_{\mathbf{k}}$ of original modes exhibits damped-harmonic envelope oscillations around its initial value. In the region of $\mathbf{k}>\mathbf{k}_{2}$ (except for $\mathbf{k}=\mathbf{k}_{3},\mathbf{k}_{4}$), the trend of $A_{\mathbf{k}}$ is similar to that of $A_{\mathbf{total}}$, indicating that collective modes with these momenta are dominant in the excitation (Fig. 4b right). Normally, in the absence of mode coupling, $A_{\mathbf{k}}$ is expected to return to near its initial value or exhibit a pronounced drop after entering the nonlinear stage. Thus we think these new collective modes are coupled each other and eventually cannot be decoupled. In the above region, the weak and peakless oscillations of $A_{\mathbf{k}}$ also reflect radial coupling among the modes. Notably, the three modes ($\mathbf{k}_{1},\mathbf{k}_{3},\mathbf{k}_{4}$) of the ground state retain distinct peaks throughout the whole dynamical evolution. Obviously these three modes have no radial coupling to other additional collective modes, leading to the decline in $A_{\mathbf{k}_{1,3,4}}$ because of the energy transfer from these modes to the remaining uncoupled modes (Fig. 4b left).
From Fig. 3(b), we observe the new modes emerge along specific directions when $\mathbf{k}\leq\mathbf{k}_{3}$. Inspired by the dynamical patterns in momentum-space, we wonder whether the newly appearing collective modes are coupled angularly. To quantify the distribution of these modes along the angular direction, we also evaluate $A_{\phi}=\int_{\phi}\tilde{n}^{\prime}(\mathbf{\mathbf{k}_{\phi}})d\mathbf{\mathbf{k}_{\phi}}$, where $\tilde{n}^{\prime}(\mathbf{\mathbf{\mathbf{k}_{\phi}}})=\int n(\mathbf{r})e^{-i\mathbf{k}_{\phi}\cdot\mathbf{r}}d\mathbf{r}/2\pi$ with magnitude $\left|\mathbf{k}_{\phi}\right|=\left|\mathbf{k}\right|$ and angle $\phi$. Consistent with our expectations, new modes emerge along the directions ($\phi_{2},\phi_{3}$) defined by $\mathbf{k}_{2}$ and $\mathbf{k}_{3}$ (Fig. 4c left), where $\phi_{2}=\frac{(2j+1)\pi}{6}$ and $\phi_{3}=\frac{j\pi}{3}$ with j=0,1,,5j=0,1,...,5. Interestingly, for $\mathbf{k}>\mathbf{k}_{3}$, $A_{\phi}$ consists of a series of distinct peaks (Fig. 4c right), which means the collective modes are angularly uncoupled.
IV.2 The real-space Moiré quasicrystal patterns
Figure 5: The dynamical evolution of real-space patterns with different modulation frequencies. (a) The patterns with $a_{m}=50a_{0}$ and $\omega=200$ Hz (top), $\omega=252$ Hz (middle) and $\omega=300$ Hz (bottom). (b) The corresponding $A_{total}$
of (a).
@@ IV.2 The real-space Moiré quasicrystal patterns
The rotational symmetry of the patterns significantly depends on the modulation amplitude $a_{m}$ and frequency $\omega$. We first fix the modulation amplitude and vary the frequency. Increasing the modulation frequency breaks the rotational symmetry of the system. For example, at $a_{m}=50a_{0}$ and low $\omega$, the pattern exhibits $D_{6}$ symmetries and other distinct $D_{6}$ patterns subsequently emerge during the third stage. Each of these non-subharmonic patterns represent a new symmetrical structure. Upon increasing $\omega$ to approximately 252 Hz, the $A_{total}$ decreases significantly, indicating only a minor DW alteration and patterns with lattice characteristics (Fig. 5(b) blue curve). The blue curve in Fig. 5(b) is notably smoother than the other two. We think this frequency matches exactly with the inherent frequency of a collective excitation mode. Under this frequency, system is excited without a preparation stage. In other words, energy is injected into the system directly and efficiently from the drive field, causing a smoother curve. This phenomenon differs from the collective mode momentum of DW yet does not conflict with it. For frequencies above this frequency, the formation of moiré quasicrystal patterns is suppressed, while the duration of the lattice patterns is extended.
@@
IV.3 The momentum-space Moiré quasicrystal patterns
We further analyze the symmetry of the DW distribution in momentum space based on Fourier analysis $\psi(\boldsymbol{k})=\int e^{-i\mathbf{k}\cdot\mathbf{r}}{\psi}(\mathbf{r})d\mathbf{r}/2\pi$. The Fourier DW $n(\boldsymbol{k})$ also exhibits high rotational symmetry, typically forming $D_{6}$ patterns. These momentum-space patterns are consistent with real-space patterns at specific times. At $a_{m}=100a_{0}$ and $\omega=200$ Hz, the momentum-space pattern forms a hexagram nearly identical to the real-space pattern at 36.2 ms, but rotated clockwise by $60^{\circ}$. Similarly, at $a_{m}=75a_{0}$ and $\omega=200$ Hz, the patterns in both spaces match well at 42, 44, and 64.5 ms (Fig. 7(b)). At 64.5 ms in particular, both patterns exhibit a symmetrical butterfly structure. The momentum-space patterns also do not exhibit subharmonicity. However, unlike the real-space patterns, some momentum-space patterns exhibit only $D_{4}$ symmetries at the center.